Problem 5-4 Calculating Interest Rates [LO3]

Solve for the unknown interest rate in each of the following (Do not round intermediate calculations. Enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.):

  

Present Value Years Interest Rate Future Value
$ 260 3      %     $ 307
380 17               1,107
41,000 18               185,586
40,261 20               342,595
Answer #1

r = (FV / PV)1/t – 1

r = ($307 / $260)1/3 – 1 = 0.0570, or 5.70%

r = ($1,107 / $380)1/17 – 1 = 0.0649, or 6.49%

r = ($185,586 / $41,000)1/18 – 1 = 0.0875, or 8.75%

r = ($342,595 / $40,261)1/20 – 1 = 0.1130, or 11.30%